# Vertically resolved minimal-set k-distribution for thermal infrared absorption in the Venus atmosphere

### v1

@@ 2.3. Resulting k-tables
The resulting set of $k$-distribution terms is summarized in Table 1. Over the longwave interval 10–6000 cm-1, we obtain $N=16$ bands and only 32 k-terms. In terms of the number of radiative transfer solves, this parameterization is more compact than existing correlated-$k$ implementations for Venus. The scheme introduced by [takahashi2023development], which adopts a very similar band layout (17 bands over 10–6425 cm-1), requires about 3.5 times more solutions of the transport equations than our scheme. Other longwave parameterizations developed for Venus GCMs, such as the Net Exchange Rate–based scheme of [eymet2009net] later implemented by [lebonnois2015analysis], employ considerably finer spectral and $k$-distribution discretizations and therefore involve a larger number of radiative transfer evaluations per call than the FKDM approach presented here.
@@
For each band, we provide — in addition to the set of $M_{i}$ k-terms — a single representative spectral point $\tilde{\nu}_{i}^{\text{cld}}$ (Table 1) to be used for the computation of cloud optical properties. In the longwave spectral range considered here, cloud extinction and single-scattering properties vary only weakly across an individual band, so we assume them to be constant within the band and evaluate them at the prescribed $\tilde{\nu}_{i}^{\text{cld}}$. Additionally, as mentioned in section 4., a band-averaged Planckian profile is introduced. Thus, each band $i$ is fully characterized by its number of k-terms $M_{i}$ and by its representative spectral point $\tilde{\nu}_{i}^{\text{cld}}$ and Planck function values, which together define the inputs required by the radiative transfer solver.
@@
For the second $k$-term in the 600–760 cm-1 band, we compared the corresponding logarithms of effective cross sections, $S=\ln\sigma_{\mathrm{eff}}$, obtained for different representative Venus temperature profiles as functions of $w=\ln p$. In the interval $w\in(-4,7)$, corresponding to altitudes $z\in(100,50)$ km, the resulting $S(w)$ profiles remain close to each other, reflecting the modest temperature differences within this altitude range. A similar behavior was found for all other spectral bands, indicating weak sensitivity of the effective cross sections to realistic variations in the Venus thermal structure. Based on that, a single ‘‘baseline’’ effective cross-section profile $S^{\ast}(w)$ can be adopted for all considered temperature profiles, except in the 600–760 and 760–900 cm-1 intervals. For these two bands we retain an explicit temperature dependence by applying, at each level $w$, an empirical correction
@@
where $T^{\ast}(w)$ is the baseline temperature profile and $\alpha(w)$ is determined from numerical experiments. In the final parameterization, the functions $S^{\ast}(w)$ and $\alpha(w)$ are provided in tabular form within the Fortran code, for more details see section 4..
Figures 2 and 3 compare upward radiative fluxes and cooling rates computed with the FKDM parameterization against the corresponding line-by-line reference calculations for representative Venus atmospheric profiles shown in Fig. 1. Very close agreement between the two solutions is evident for the upward fluxes. For the cooling rates the error does not exceed $\sim$1 K/day for all altitude ranges, as the ability to finely control the accuracy is partially constrained by numerical effects associated with the chosen vertical grid resolution, which was selected to keep cloud optical properties constant within each model layer.
@@
3. Radiative transfer model
@@
We employed a reference line-by-line radiative transfer model based on the Monte-Carlo method [liou2002introduction] and its algorithmic implementation developed in [fomin1998model, fomin2006monte]. The same code family has been used previously in our FKDM $k$-distribution study for the UV region [Fomin2022], and the Earth version has participated in radiation-code intercomparisons for climate applications [Oreopoulos2012continual]. The model provides an accurate treatment of gaseous absorption and particulate multiple scattering for longwave radiation in plane-parallel slab atmospheres with one or several cloud and aerosol layers. Multiple scattering is simulated explicitly, and the residual error in fluxes and cooling rates decreases as $K^{-1/2}$ with the total number of photon histories $K$, independently of the number of spectral points [fomin2006monte]. This statistical convergence, together with the fact that the error is governed by photon statistics rather than by the spectral grid, makes the approach particularly suitable for broad-band longwave calculations in highly scattering, optically thick cloud layers [fomin2006monte]. Consequently, in the present work we simply adopt the native high-resolution grid at which monochromatic absorption coefficients are precomputed, with a spacing of $1/2048\penalty 10000\ \mathrm{cm}^{-1}$, letting the Monte-Carlo statistics control the overall accuracy of the reference solution.
We present the resulting $k$-distribution set for the altitude range of 0-100 km. Choosing this altitude range largely depends upon the altitude limit for reliable reference modeling with existing thermal profiles. Temperature structure in 100-150 km region remains observationally underconstrained, while obtained temperature change rates can have substantial uncertainties even in 90-100 km range as mentioned by [Haus2015, Haus2016]. Only three gaseous species are treated as main absorbers in the modeling atmosphere: CO2, H2O, and SO2. Trace gases (e.g. CO, OCS, HCl, HF) are omitted from the generation of the $k$-distribution, as their primary importance lies in spectroscopic details in narrow windows rather than in energetic studies [Haus2015, Haus2016] or in applications for GCMs [mendonca2015new].
@@
3.1. Atmospheric profiles
[…]
@@ 3.1. Atmospheric profiles
Initial VIRA-like data provide only total column abundance, so to decouple individual densities, volume mixing rations of each species must be introduced. The volume mixing ratio of CO2 is assumed to be constant with altitude and is fixed at 0.965. H2O abundance is set to 32.5 ppmv below 50 km, consistent with near-IR determinations of $\sim$30–35 ppmv in the lower atmosphere [Arney2014, Fedorova2016], and to 3 ppmv near and above the cloud tops ($\sim$70 km), consistent with nadir and occultation measurements of $\sim$1–3 ppmv [Cottini2012, Fedorova2016]. SO2 abundances are taken from Fig.5 of model input by [Haus2015], that describe a vertically uniform lower-atmosphere mixture below 50 km, with volume mixing ratios of 32.5 and 150 ppmv, respectively.
3.2. Absorption by gaseous components
@@
Our recently introduced MARFA code [Razumovskiy2025MARFA] is employed for calculation of absorption coefficients based on line parameters. MARFA computes absorption coefficients at a high spectral resolution of $1/2048\penalty 10000\ \mathrm{cm^{-1}}$, which is maintained into the far wings of spectral lines through the use of an efficient multi-grid interpolation algorithm [fomin1995effective]. Particular emphasis is placed on evaluating line shapes with large cut-offs, an essential feature in presence of uncertainties in spectroscopic parameters and continuum absorption in Venus modeling. In the absence of a dedicated, fully consistent Venus spectroscopic database, line parameters must be assembled from several sources. For spectroscopic input, we utilized the HITEMP database [rothman2010HITEMP] for CO2 and H2O, which provides comprehensive high-temperature line lists essential for Venus’ broad thermal range. For SO2, we relied on HITRAN2020 database [Gordon2022].
@@
The choice of continuum absorption parameterization in Venus radiative transfer modeling depends on the intended application. Pressure-induced continuum has a non-negligible influence in retrieval studies, for example on cloud properties or minor-species abundances [Haus2013selfconsistent], or in simulations of deep-atmosphere net fluxes for comparison with entry probe measurements [lee2016sensitivity]. In such studies—typically restricted to relatively narrow spectral intervals—continuum absorption is often described either by physically based laboratory determinations [moskalenko1979pressure, gruszka1997roto, richard2012new, stefani2013experimental, tran2024collision] or by introducing additional empirical opacity to match observations in near-infrared windows [bezard1990chi-factor, de1995water, marcq2006remote, Bezard2011]. A sensitivity analysis by [Haus2015] demonstrated that, for the purpose of computing broadband net fluxes and heating rates, continuum absorption can be safely neglected; this conclusion was further supported by their energy-balance study [Haus2016]. In those studies, the relevant collisional line-mixing and pressure-induced effects were incorporated implicitly through sub-Lorentzian $\chi$-factors and far-wing truncation. Following the recommendations of [Haus2015, Haus2016], we similarly applied the sub-Lorentzian correction of [Tonkov1996] to the CO2 far wings and truncated lines at 250 cm-1. No intensity cut-off was imposed, as MARFA efficiently treats the full line list. Because the primary purpose of our $k$-distribution tables is to enhance the computational efficiency of radiative transfer modules in Venus climate modeling (where emphasis is made on broad spectral ranges), adopting a Haus-type treatment of continuum absorption [Haus2015, Haus2016] is well justified. In future, our line-by-line Monte-Carlo model will be released as a standalone tool suitable for a broader class of radiative transfer problems and we plan to incorporate a physically based continuum absorption model.
@@
In a recent study, [chaverot2025spect] introduced updated $\chi$-factor parameterizations for several planetary environments, including Venus, based on a comprehensive compilation and re-analysis of laboratory data. In their framework, the definition of continuum absorption is made more rigorous by adopting a fixed line-wing cut-off at 25 cm-1 from the line center. The same $\chi$-factor formalism is then used to construct separate CO2 line-wing continua over 25–1500 cm-1, as well as dedicated pressure-induced continuum files. In future versions of our $k$-distribution scheme (e.g. for short-wave radiation) we plan to adopt these new $\chi$-factors and continuum absorption parameterizations.
3.3. Cloud aerosol properties: vertical structure and microphysics
@@
Cloud particle concentrations are adopted from the initial cloud model introduced by [Haus2015], which is based on the VIRTIS-derived self-consistent retrievals from [Haus2013selfconsistent]. Per-mode concentrations follow the analytical expressions provided in Table 1 of their study [Haus2015]. Log-normal size distribution parameters are likewise drawn from [Pollack1993]. The model incorporates four modes of spherical liquid droplets, each composed of a sulfuric acid-water mixture (H2SO4-H2O) with a uniform concentration of 75 wt% H2SO4 (by weight) across all modes and altitudes. Refractive index data are sourced from the study by [Palmer1975]. Based on Mie scattering theory [wiscombe1980improved], we employed the same computational code as in our prior work [Fomin2022] to derive the cloud microphysical optical properties. These parameters serve as inputs for our radiative transfer model. Over the 10–6000 cm-1 spectral interval considered in the present study, cloud droplets behave as strong absorbers throughout most of the thermal infrared ($\lesssim 3500$ cm-1), approaching nearly conservative scattering only in the short-wave near-IR tail ($\gtrsim 4000$ cm-1) [Haus2015].
4. Code and data
@@
To support the use of the longwave FKDM parameterization in radiative transfer models and general circulation models, we provide a compact Fortran implementation in a public repository [^1] https://github.com/Razumovskyy/Venus-longwave-fkdm. The repository contains a Fortran driver which computes vertical profiles of $k$-distribution terms and band-integrated Planck functions for arbitrary atmospheric columns. The code is organized as a Fortran Package Manager (fpm) project, such that compilation and execution require only fpm run. By default, two diagnostic files are produced: VAC, containing the volume absorption coefficients $k_{i,j}(z)$, and PLANCK, containing the corresponding band-integrated Planck functions. These reproduce the quantities a host GCM or radiative transfer solver would typically obtain through direct calls to the FKDM routines. The implementation is modular and designed to operate either as a stand-alone tool for one-dimensional radiative transfer tests or as transferable module component for any Venus GCM. The repository is equipped with a Readme file to facilitate onboarding.
@@
A key advantage of this design is that users do not interact with the line-by-line or Monte Carlo calculations, nor with the reference atmospheres employed during construction of the $k$-distributions. The FKDM module essentially acts as a lightweight mapping:
$\bigl(P(z),\,T(z),\,q_{\mathrm{CO_{2}}}(z),\,q_{\mathrm{H_{2}O}}(z),\,q_{\mathrm{SO_{2}}}(z)\bigr)\quad\longrightarrow\quad\{k_{i,j}(z),\,B_{i}(z)\},$
@@
5. Conclusion
We have developed a minimal-set $k$-distribution parameterization for longwave gaseous absorption in the Venus atmosphere and validated it against a high-resolution line-by-line Monte-Carlo radiative transfer model. The FKDM approach produces only 32 spectral channels across 16 bands for CO2, H2O and SO2, while maintaining accuracy comparable to standard correlated-$k$ implementations and potentially beyond that. The method natively accounts for vertical inhomogeneity, avoids assumptions about inter-level correlation of monochromatic absorption, and allows direct control over the accuracy–efficiency balance during construction of the k-terms.
Validation across a set of representative Venus temperature profiles demonstrates that upward radiative fluxes computed with the FKDM scheme closely reproduce the corresponding line-by-line results up to $\sim$100 km, reflecting the described exact flux-matching strategy employed during the construction of the $k$-distribution terms. Cooling rates are reproduced with good accuracy throughout the lower and middle atmosphere, remaining consistent with line-by-line calculations up to $\sim$90 km. The compactness of the resulting parameterization yields a substantial reduction in radiative transfer evaluations, making the FKDM approach well suited for Venus GCMs and radiative–convective models.

### v2

@@ 2.3 Resulting k-table
The resulting set of $k$-distribution terms is summarized in Table 1. Over the longwave interval 10–6000 cm-1, we obtain $N=16$ bands and only 32 k-terms. In terms of the number of radiative transfer solves, this parameterization is more compact than existing correlated-$k$ implementations for Venus. The scheme introduced by takahashi2023development, which adopts a very similar band layout (17 bands over 10–6425 cm-1), requires about 3.5 times more solutions of the transport equations than our scheme. Other longwave parameterizations developed for Venus GCMs, such as the Net Exchange Rate–based scheme of eymet2009net later implemented by lebonnois2015analysis, employ considerably finer spectral and $k$-distribution discretizations and therefore involve a larger number of radiative transfer evaluations per call than the FKDM approach presented here.
@@
For each band, we provide — in addition to the set of $M_{i}$ k-terms — a single representative spectral point $\tilde{\nu}_{i}^{\text{cld}}$ (Table 1) to be used for the computation of cloud optical properties. In the longwave spectral range considered here, cloud extinction and single-scattering properties vary only weakly across an individual band, so we assume them to be constant within the band and evaluate them at the prescribed $\tilde{\nu}_{i}^{\text{cld}}$. Additionally, as mentioned in section 4, a band-averaged Planckian profile is introduced. Thus, each band $i$ is fully characterized by its number of k-terms $M_{i}$ and by its representative spectral point $\tilde{\nu}_{i}^{\text{cld}}$ and Planck function values, which together define the inputs required by the radiative transfer solver.
@@
For the second $k$-term in the 600–760 cm-1 band, we compared the corresponding logarithms of effective cross sections, $S=\ln\sigma_{\mathrm{eff}}$, obtained for different representative Venus temperature profiles as functions of $w=\ln p$. In the interval $w\in(-4,7)$, corresponding to altitudes $z\in(100,50)$ km, the resulting $S(w)$ profiles remain close to each other, reflecting the modest temperature differences within this altitude range. A similar behavior was found for all other bands, indicating weak sensitivity of the effective cross sections to realistic variations in the Venus thermal structure. Based on that, a single ‘‘baseline’’ effective cross-section profile $S^{\ast}(w)$ can be adopted for all considered temperature profiles, except in the 600–760 and 760–900 cm-1 intervals. For these two bands we retain an explicit temperature dependence by applying, at each level $w$, an empirical correction
@@
where $T^{\ast}(w)$ is the baseline temperature profile and $\alpha(w)$ is determined from numerical experiments. In the final parameterization, the functions $S^{\ast}(w)$ and $\alpha(w)$ are provided in tabular form within the Fortran code and are accessed internally by the FKDM module described in section 4.
Figures 2 and 3 compare upward radiative fluxes and cooling rates computed with the FKDM parameterization against the corresponding line-by-line reference calculations for representative Venus atmospheric profiles shown in Fig. 1. Close agreement between the two solutions is evident for the upward fluxes and is of <2% below 95 km. As a basic consistency check, the globally averaged outgoing thermal flux at TOA is expected to be $157\pm 6\penalty 10000\ \mathrm{W\,m^{-2}}$ (titov2007radiation; eymet2009net), corresponding to about 3.8 % of plausible deviation. The sensitivity study by Haus2015 reports that uncertainties in cooling rates below 95 km are about 0.25 $\mathrm{K\,day^{-1}}$. In our implementation, discrepancies do not exceed approximately 1.5 $\mathrm{K\,day^{-1}}$ below 95 km, decreasing to below 1.2 $\mathrm{K\,day^{-1}}$ below 90 km and to 0.3 $\mathrm{K\,day^{-1}}$ below 75 km approaching the level of uncertainties inherent to reference calculations.
@@
3 Radiative transfer model
@@
We employed a reference line-by-line radiative transfer model based on the Monte-Carlo method (liou2002introduction) and its algorithmic implementation developed in fomin1998model; fomin2006monte. The same code family has been used previously in our FKDM $k$-distribution study for the UV region (Fomin2022), and the Earth version has participated in radiation-code intercomparisons for climate applications (Oreopoulos2012continual). The model provides an accurate treatment of gaseous absorption and particulate multiple scattering for longwave radiation in plane-parallel slab atmospheres with one or several cloud and aerosol layers. Multiple scattering is simulated explicitly, and the residual error in fluxes and cooling rates decreases as $K^{-1/2}$ with the total number of photon histories $K$, independently of the number of spectral points (fomin2006monte). This statistical convergence, together with the fact that the error is governed by photon statistics rather than by the spectral grid, makes the approach particularly suitable for broadband longwave calculations in highly scattering, optically thick cloud layers (fomin2006monte). Consequently, in the present work we simply adopt the native high-resolution grid at which monochromatic absorption coefficients are precomputed, with a spacing of $1/2048\penalty 10000\ \mathrm{cm}^{-1}$, letting the Monte-Carlo statistics control the overall accuracy of the reference solution.
We present the resulting $k$-distribution set for the altitude range of 0-100 km. Choosing this altitude range largely depends upon the altitude limit for reliable reference modeling with existing thermal profiles. Temperature structure in 100-150 km region remains observationally underconstrained, while obtained temperature change rates can have substantial uncertainties even in 90-100 km range as mentioned by Haus2015; Haus2016. Only three gaseous species are treated as main absorbers in the modeling atmosphere: CO2, H2O, and SO2. Trace gases (e.g. CO, OCS, HCl, HF) are omitted from the generation of the $k$-distribution, as their primary importance lies in spectroscopic details in narrow windows rather than in energetic studies (Haus2015; Haus2016) or in applications for GCMs (mendonca2015new).
@@
3.1 Atmospheric profiles
We utilized a set of three representative and realistically contrasting atmospheric profiles: the low-latitude midnight profile introduced by Seiff1983, VIRA-1 midnight 75 latitude profile from Seiff1985 and a combined profile with VIRA-2 75 latitude day-side profile introduced by Zasova2006 for solar longitude sector $L_{s}=20^{\circ}-90^{\circ}$ in the middle atmosphere 50-100 km and relevant VIRA-1 profile below 50 km. Temperature profiles are depicted on Fig. 1.
@@ 3.1 Atmospheric profiles
Initial VIRA-like data provide only total column abundance, so to decouple individual densities, volume mixing rations of each species must be introduced. The volume mixing ratio of CO2 is assumed to be constant with altitude and is fixed at 0.965. H2O abundance is set to 32.5 ppmv below 50 km, consistent with near-IR determinations of $\sim$30–35 ppmv in the lower atmosphere (Arney2014; Fedorova2016), and to 3 ppmv near and above the cloud tops ($\sim$70 km), consistent with nadir and occultation measurements of $\sim$1–3 ppmv (Cottini2012; Fedorova2016). SO2 abundances are taken from Fig.5 of model input by Haus2015, that describe a vertically uniform lower-atmosphere mixture below 50 km, with volume mixing ratios of 32.5 and 150 ppmv, respectively.
3.2 Absorption by gaseous components
@@
Our recently introduced MARFA code (Razumovskiy2025MARFA) is employed for calculation of absorption coefficients based on line parameters. MARFA computes absorption coefficients at a high spectral resolution of $1/2048\penalty 10000\ \mathrm{cm^{-1}}$, which is maintained into the far wings of spectral lines through the use of an efficient multi-grid interpolation algorithm (fomin1995effective). Particular emphasis is placed on evaluating line shapes with large cut-offs, an essential feature in presence of uncertainties in spectroscopic parameters and continuum absorption in Venus modeling. In the absence of a dedicated, fully consistent Venus spectroscopic database, line parameters must be assembled from several sources. For spectroscopic input, we utilized the HITEMP database (rothman2010HITEMP) for CO2 and H2O, which provides comprehensive high-temperature line lists essential for Venus’ broad thermal range. For SO2, we relied on HITRAN2020 database (Gordon2022).
@@
The choice of continuum absorption parameterization in Venus radiative transfer modeling depends on the intended application. Pressure-induced continuum has a non-negligible influence in retrieval studies, for example on cloud opacity and related cloud parameters (Haus2013selfconsistent; evdokimova2025cloud), or in simulations of deep-atmosphere net fluxes for comparison with entry probe measurements (lee2016sensitivity). In such studies—typically restricted to relatively narrow spectral intervals—continuum absorption is often described either by physically based laboratory determinations (moskalenko1979pressure; gruszka1997roto; richard2012new; stefani2013experimental; tran2024collision) or by introducing additional empirical opacity to match observations in near-infrared windows (bezard1990chi-factor; de1995water; marcq2006remote; Bezard2011). A sensitivity analysis by Haus2015 demonstrated that, for the purpose of computing broadband net fluxes and heating rates, continuum absorption can be safely neglected; this conclusion was further supported by their energy-balance study (Haus2016). In those studies, the relevant collisional line-mixing and pressure-induced effects were incorporated implicitly through sub-Lorentzian $\chi$-factors and far-wing truncation. Following the recommendations of Haus2015; Haus2016, we similarly applied the sub-Lorentzian correction of Tonkov1996 to the CO2 far wings and truncated lines at 250 cm-1. No intensity cut-off was imposed, as MARFA efficiently treats the full line list. Because the primary purpose of our $k$-distribution tables is to enhance the computational efficiency of radiative transfer modules in Venus climate modeling (where emphasis is made on broad spectral ranges), adopting a Haus-type treatment of continuum absorption (Haus2015; Haus2016) is well justified. In future, our line-by-line Monte-Carlo model will be released as a standalone tool suitable for a broader class of radiative transfer problems and we plan to incorporate a physically based continuum absorption model.
@@
In a recent study, chaverot2025spect introduced updated $\chi$-factor parameterizations for several planetary environments, including Venus, based on a comprehensive compilation and re-analysis of laboratory data. In their framework, the definition of continuum absorption is made more rigorous by adopting a fixed line-wing cut-off at 25 cm-1 from the line center. The same $\chi$-factor formalism is then used to construct separate CO2 line-wing continua over 25–1500 cm-1. Additionally, dedicated continuum files, accounting for collision-induced and dimer absorption, are provided. In future versions of our $k$-distribution scheme (e.g. for short-wave radiation) we plan to adopt these new $\chi$-factors and continuum absorption parameterizations.
3.3 Cloud aerosol properties
@@
Cloud particle concentrations are adopted from the initial cloud model introduced by Haus2015, which is based on the VIRTIS-derived self-consistent retrievals from Haus2013selfconsistent. Per-mode concentrations follow the analytical expressions provided in Table 1 of their study (Haus2015). Log-normal size distribution parameters are likewise drawn from Pollack1993. The model incorporates four modes of spherical liquid droplets, each composed of a sulfuric acid-water mixture (H2SO4-H2O) with a uniform concentration of 75 wt% H2SO4 (by weight) across all modes and altitudes. Refractive index data are sourced from the study by Palmer1975. Based on Mie scattering theory (wiscombe1980improved), we employed the same computational code as in our prior work (Fomin2022) to derive the cloud microphysical optical properties. These parameters serve as inputs for our radiative transfer model. Over the 10–6000 cm-1 spectral interval considered in the present study, cloud droplets behave as strong absorbers throughout most of the thermal infrared ($\lesssim 3500$ cm-1), approaching nearly conservative scattering only in the short-wave near-IR tail ($\gtrsim 4000$ cm-1) (Haus2015).
4 Code and data
@@
To support the use of the longwave FKDM parameterization in radiative transfer models and general circulation models, we provide a compact Fortran implementation in a public repository [^1] https://github.com/Razumovskyy/Venus-longwave-fkdm. The repository contains a Fortran driver which computes vertical profiles of $k$-distribution terms and band-integrated Planck functions for arbitrary atmospheric columns. The code is organized as a Fortran Package Manager (fpm) project, such that compilation and execution require only fpm run. By default, two diagnostic outputs are produced: (i), containing the volume absorption coefficients $k_{i,j}(z)$, and (ii), containing the corresponding band-integrated Planck functions. These reproduce the quantities a host GCM or radiative transfer solver would typically obtain through direct calls to the FKDM routines. The implementation is modular and designed to operate either as a stand-alone tool for one-dimensional radiative transfer tests or as transferable module component for any Venus GCM. The repository is equipped with a Readme file to facilitate onboarding.
@@
From the user’s perspective the FKDM module essentially acts as a black box and provides a lightweight mapping:
$\bigl(P(z),\,T(z),\,q_{\mathrm{CO_{2}}}(z),\,q_{\mathrm{H_{2}O}}(z),\,q_{\mathrm{SO_{2}}}(z)\bigr)$
$\longrightarrow\quad\{k_{i,j}(z),\,B_{i}(z)\},$
@@
5 Conclusion
We have developed a minimal-set $k$-distribution parameterization for longwave gaseous absorption in the Venus atmosphere and validated it against a high-resolution line-by-line Monte-Carlo radiative transfer model. The FKDM approach produces only 32 spectral channels across 16 bands for CO2, H2O and SO2, while maintaining accuracy comparable to standard correlated-$k$ implementations and potentially beyond that. The method natively accounts for vertical inhomogeneity, avoids assumptions about inter-level correlation of monochromatic absorption, and allows direct control over the accuracy–efficiency balance during construction of the k-terms. Although developed and validated here for Venus case, the method is general and can be adapted to other planetary regimes.
Validation across a set of representative Venus temperature profiles demonstrates that upward radiative fluxes computed with the FKDM scheme closely reproduce the corresponding line-by-line results up to 95 km, reflecting the flux-matching strategy employed during the construction of the $k$-distribution terms. Cooling rates are reproduced with acceptable accuracy throughout the lower and middle atmosphere, remaining consistent with line-by-line calculations up to 90 km. The compactness of the resulting parameterization yields a substantial reduction in radiative transfer evaluations, making the FKDM approach well suited for Venus GCMs and radiative–convective models.